Jeffery currently owes $8,000 to a friend who is charging him interest of 2.80% p.m. He is required to settle the loan with two equal payments, one today and the other in seven months. Calculate the size of the payments using secen months as the focal date. Round to the nearest cent.
Answer #1

Let 'X' be the payment today.

Then payment after 7 months = ( 8000 - X )( 1+ 0.028 )^7

= ( 8000- X ) 1.2133

Since payment now and payment after seven months are equal, which means

X = ( 8000 - X ) 1.2133

X = 9706.4 - 1.2133X

2.2133 X = 9706.4

X= 9706.4 / 2.2133 = 4385.49

Therefore Jeffery has to pay $ 4385.49 today and after 7 months